OpenStudy (anonymous):

A 32-foot ladder leans against a building so that the top of the ladder touches the building at a point 23 feet above the ground. To the nearest tenth of a foot, how far from the bottom of the building is the bottom of the ladder? Please help! I'll medal!

OpenStudy (ageta):

Use the Pythagorean theorem for this. a^2+b^2=c^2 The ladder is c The wall is b The ground is a Rearrange the equation so that you can solve for a. a^2=c^2-b^2 Substitute values. a^2=32^2+23^2 Simplify a^2=1024-529 a^2=495 Square both sides. √a^2=√495=22.2 Answer: 22.2

OpenStudy (ageta):

Answer: 22.2

OpenStudy (ageta):

what do you guys think?

OpenStudy (xapproachesinfinity):

seems fine with me

OpenStudy (anonymous):

Thank you!

OpenStudy (ageta):

you welcome